OpenStudy (anonymous):

[9.02] What are the x-intercept(s) of the graph of y − x^2 = −8x − 9?

OpenStudy (anonymous):

add x^2 to the right y=x^2-8x-9 factor and set each equation to zero

OpenStudy (anonymous):

(x+1)(X-9)=0 x=-1 and x=9